# Towards Geo-Culturally Grounded LLM Generations

### v2

@@ Abstract
Generative large language models (LLMs) have been demonstrated to have gaps in diverse, cultural knowledge across the globe. We investigate the effect of retrieval augmented generation and search-grounding techniques on the ability of LLMs to display familiarity with a diverse range of national cultures. Specifically, we compare the performance of standard LLMs, LLMs augmented with retrievals from a bespoke knowledge base (i.e., KB grounding), and LLMs augmented with retrievals from a web search (i.e., search grounding) on a series of cultural familiarity benchmarks. We find that search grounding significantly improves the LLM performance on multiple-choice benchmarks that test propositional knowledge (e.g., the norms, artifacts, and institutions of national cultures), while KB grounding’s effectiveness is limited by inadequate knowledge base coverage and a suboptimal retriever. However, search grounding also increases the risk of stereotypical judgments by language models, while failing to improve evaluators’ judgments of cultural familiarity in a human evaluation with adequate statistical power. These results highlight the distinction between propositional knowledge about a culture and open-ended cultural fluency when it comes to evaluating the cultural familiarity of generative LLMs.
@@ 1 Introduction
Contemporary large language models (LLMs) are pre-trained on huge corpora of natural language text Radford et al. (2019) and then fine-tuned using human feedback to improve their quality Bai et al. (2022). During both processes, it is possible for text from a particular culture or cultures to be over-represented in the training data Dodge et al. (2021) and for the perspectives, norms, and mores of specific cultures to be over-represented in the feedback from human evaluators Prabhakaran et al. (2022); Atari et al. (2023). Consequently, there is growing recognition of the shortcomings of generative LLMs with respect to their ability to represent and serve people of diverse geo-cultural backgrounds at the global scale Adilazuarda et al. (2024); Pawar et al. (2024); Agarwal et al. (2025). Models tend to stereotype different cultures Jha et al. (2023); Bhutani et al. (2024), erase and simplify their representation Qadri et al. (2025), and provide very limited knowledge and context about artifacts and norms that are salient to them Myung et al. (2024); Rao et al. (2024). Despite these gaps, strategies for intervening and eliciting culturally appropriate content from the models remain under-explored.
@@
We study two strategies to improve cultural awareness of LLM generation using external knowledge. In the first strategy, we construct a bespoke cultural knowledge base (KB) and apply a retrieval augmented generation (RAG) technique Lewis et al. (2020); Gao et al. (2023) so that the input to the language model includes relevant cultural text from the knowledge base for better generation. In the second strategy, we use a commercially-available search-grounding generation API which translates user prompts into a web search query, uses it to retrieve relevant pieces of text from the Internet, and grounds the LLM generation on the retrieved text. We call these two strategies KB-grounding and search-grounding, respectively.
@@
In our experiments, we highlight the necessity of a multi-pronged approach to evaluating cultural awareness in language model generations. Specifically, we leveraged multiple benchmarks Myung et al. (2024); Rao et al. (2024); Bhutani et al. (2024) to evaluate cultural knowledge and the ability to avoid cultural stereotyping of the models equipped with the two strategies, compared with the vanilla generation baseline. We also conducted a human evaluation of open-ended model responses to various prompts designed to test cultural awareness, wherein evaluators from a specific national culture rated how well the model’s output reflected a culturally informed perspective. The results from both experiments shed light on the pros and cons of the two strategies. Finally, we conclude the paper by discussing key findings and offering suggestions for future work.
2 Improving Cultural Awareness with External Knowledge
@@
Retrieval augmented generation (RAG) is a technique for enhancing the quality of large language model generation. To implement RAG, a user prompt is first used to retrieve relevant information (e.g., text or documents) from a database, which is then added to the prompt before being passed to a generative LLM to produce a grounded response Lewis et al. (2020); Gao et al. (2023). RAG has shown to be effective in several applications, particularly those involving tasks the base LLM was not well-trained for such as fact verification Asai et al. (2023); Singal et al. (2024); Khaliq et al. (2024) and domain-specific question answering (QA) Seo et al. (2024); Xiong et al. (2024); Kim and Min (2024). Some existing work also creates bespoke knowledge bases for use with RAG to tailor outputs to their specific applications Sun et al. (2025); Li et al. (2024). Meanwhile, instead of querying a KB, other techniques retrieve external information by searching the internet, which allows access to up-to-date knowledge with high-quality ranking outputs Fan et al. (2024); Shuster et al. (2022); Lazaridou et al. (2022); Yao et al. (2022); Nakano et al. (2021); Komeili et al. (2022). In light of these successes, we study both KB-grounding and search-grounding as techniques for improving the cultural awareness of LM generations.
2.1 The Knowledge Base Grounding Strategy
We compiled culturally salient data from four large sources—CultureAtlas Fung et al. (2024), Cube Kannen et al. (2024), CultureBank Shi et al. (2024), and SeeGULL Jha et al. (2023)—to serve as our knowledge base. For each entry in each source, we converted it into into text (if it was not already), embedded it into a vector, and added it to a vector store for querying.
Fig. 1 (top) shows how the KB-grounding strategy handles an incoming prompt. First, a query rewriter extracts the important parts of the prompt to form a query. We can configure this step to test different KB queries (e.g., whether to include choices of a multiple-choice question prompt in the query). Next, we use the query to retrieve $n$ documents from the KB. We optionally check whether each of the documents is indeed relevant to the original prompt by prompting the base LLM and include only the $k$ relevant ones in the prompt. We call the process with the relevancy check step selective RAG, while non-selective RAG includes all $n$ documents in the prompt. Then, we feed the augmented prompt to the LLM to get a raw answer. For multiple-choice question answering (QA) tasks, we create a manual verbalizer to map the raw answer to one of the choices so that we can compute the model accuracy. In contrast, we use the raw answer as the final output for an open-ended generation task. See the Appendix for details about the KB, the queries, the prompt templates, and the implementation.
In Section 3, we apply KB-grounding to three different LLMs: gemini-flash-1.5 (Gemini) Team et al. (2024), gpt-4o-mini (GPT) OpenAI et al. (2024) and olmo2-7b (OLMo) OLMo et al. (2025).
2.2 The Search Grounding Strategy
Fig 1 (bottom) outlines how the search-grounding strategy works. We feed the original prompt to the search-grounding generation API which converts the prompt into a search query, inputs it to a search engine to obtain relevant text from the web pages returned by the search, integrates relevant text into the prompt, and then generates a response using an LLM based on the augmented prompt. Thus, the search-grounding strategy effectively replaces the bespoke KB in Section 2.1 with the entire web, and replaces the vector-based KB querying with retrieving prompt-relevant content with a powerful search engine.
We implemented search grounding on Google Vertex AI. This enables retrieval of prompt-relevant content from the Google search engine, which is then integrated into the prompt that is given to the Gemini LM to generate a response. Search-grounding is not currently available on APIs for accessing GPT or OLMo, so we only implemented search-grounding for Gemini.
3 Cultural Competence Benchmarks
We evaluated the KB-grounding and search-grounding strategies using multiple-choice cultural QA benchmarks.
3.1 Cultural Knowledge
Setup.
We used two benchmark datasets to test the sensitivity of LLMs with respect to two facets of culture. The first, BLEnD, contains questions about everyday cultural knowledge (such as food, sports, family, and education) in different countries Myung et al. (2024). We used the ~34k English questions from BLEnD that are related to ten countries [^1] China, Ethiopia, Greece, Indonesia, Iran, Mexico, South Korea, Spain, the United Kingdom, and the United States. represented in our bespoke KB. The second, NormAd, focuses on cultural norms and values Rao et al. (2024). Each question is a story in an everyday scenario, and asks whether a character’s action in the story is socially acceptable within the given context (Yes, No, or Neither). We experimented with two types of contexts – Country and Country+Value. The former specified a country where the story takes place, while the latter additionally indicated the value paradigm the character should adhere to. NormAd has ~2.6k questions in total.
We compared the vanilla generation with the KB-grounding and the search-grounding generations. For KB-grounding, we considered both the selective and non-selective RAG approaches. For BLEnD specifically, because answer choices are potentially relevant to the semantic content of the question, we considered query rewriting both “with choices” and “without choices” included in the KB query.
Results.
Fig. 2 (left) presents the results of the BLEnD benchmark. A repeated-measures ANOVA revealed a significant effect of strategy on accuracy across all three LLMs. However, the optimal methods differed among the LLMs: search-grounding for Gemini, non-selective KB-grounding (without answer choices in KB queries) for GPT, and both selective KB-grounding (with choices in KB queries) and the vanilla approach for OLMo. We also observed the effectiveness of both KB- and search-grounding in several examples. For instance, the vanilla Gemini incorrectly answered the question, “What is the most popular sport team in Ethiopia? (A) coffee (B) lg twins (C) persepolis (D) real madrid.” It selected (D), likely due to Real Madrid’s high frequency in the training corpora concerning popular sport teams. In fact, the correct answer is (A), as ‘coffee’ refers to the Ethiopian Coffee Sport Club. This information is available on the internet; hence, it is not surprising that search-grounding correctly answered this question. Notably, search-grounded Gemini achieves 74.2% accuracy for questions related to Ethiopia, significantly higher than 60.3% of the vanilla baseline and 62.9% of the best KB grounding setting. Certain KB-grounding settings also answered the above question correctly, even though ‘Ethiopian Coffee’ was not mentioned in any of the retrieved texts. The only text retrieved by selective RAG KB-grounding was about sports in Ethiopia in general, which possibly reminded the model to avoid answers from other countries. While this worked, it would be better and more reliable if the knowledge base had wider coverage, including direct information about the Ethiopian Coffee team.
The middle two charts in Fig. 2 present the NormAd benchmark results for both Country and County+Value context types. Consistent with BLEnD, search-grounded Gemini performed best for both context types compared to other strategies. For GPT, selective KB-grounding outperformed alternatives in the Country context, while the vanilla strategy performed best in the Country+Value context. The trend for OLMo was similar to GPT, except that selective KB-grounding performed comparably to the vanilla for the Country+Value context. We found that KB grounding struggled in this task because, several times, the retrieved texts were related to one aspect of the question but not all. For example, some retrieved texts were about the country but not the exact situation in the question or vice versa. This could cause the LLMs to give wrong answers. The relevancy check step implemented in selective RAG mitigated this issue, albeit to a limited extent. We posit that a better tuned KB retriever would further address this problem (only if the pertinent fact is present in the KB Joren et al. (2024)).
3.2 Stereotype Avoidance
Setup.
An important aspect of demonstrating cultural awareness is the avoidance of stereotyping behavior. Hence, we adopted the SeeGULL stereotype avoidance test proposed by Bhutani et al. (2024) to evaluate the studied strategies. Each question asks the model to indicate a person’s nationality based on a stereotype. For example,
In every question, only one country in the choices is stereotypically associated with the question. However, we considered only ‘(?)’ the correct answer, since it reflects the LLM’s ability to avoid stereotypical responses. The benchmark contains ~4.6k questions. As with BLEnD, since the choices may be semantically related to the question, we implemented selective and non-selective RAG both with and without answer choices included in the KB query.
Results.
4 Human Evaluation
Setup.
To evaluate the strategies on a more open-ended text generation task, we translated five questions from BLEnD and five questions from NormAd into open-ended prompts asking the model to tell a story set in a particular country. We adapted each of these ten questions for the ten national cultures used in the BLEnD benchmark evaluation, leaving us with 100 (country, prompt) pairs. Next, we generated responses from Gemini using the vanilla, the selective KB-grounding, and the search-grounding. For each strategy, we generated three unique responses with the temperature of 0.5. Then we recruited nine evaluators from each of the ten studied cultures to rate, on a scale from 0 to 4, how culturally informed each response was and to provide a brief justification of their score (see Appendices A.6A.7 for details).
Results.
[…]
5 Discussion and Conclusion
KB Grounding vs Search Grounding.
The advantages of search-grounding on BLEnD and NormAd speak to the vast space of cultural facts on the internet. Even the large KB we compiled here still lacked many culturally relevant facts. We also observed some bias in each knowledge source used: approximately 19% of CultureAtlas entries and 25% of CultureBank entries concern the culture of the United States. While the web as a whole remains biased towards Western sources and values Johnson et al. (2022), it is more likely to contain necessary cultural information due to its sheer scale. However, the poor performance of search-grounding on the stereotype avoidance benchmark reminds us that the context retrieved via web search could reinforce the (typically false) notion that stereotypes are factual and encourage models to affirm those stereotypes. This suggests that search-grounding is yet not a panacea for improving the cultural sensitivity of LLMs.
Knowledge vs Fluency.
The results of the human evaluation do not show that search-grounding or KB-grounding improved the cultural awareness of LLM outputs. The divergence in the performance of both strategies on the human evaluation and the multiple-choice QA benchmarks suggests that we ought to draw a distinction between two varieties of cultural awareness. On the one hand, there is a sense of cultural awareness that involves possessing propositional knowledge about a culture (i.e., knowing facts about that culture), as measured by the multiple-choice QA benchmarks, where the grounding strategies can improve the model performance. However, another sense of cultural awareness involves writing and speaking like someone with first-hand experience of and immersion in a culture, i.e., a sense of cultural fluency, as measured by our human evaluation, which revealed that the two grounding strategies was of limited value. We leave it to future work to develop strategies for improving the cultural awareness of generative language models along this second axis.
Limitations
We acknowledge the following limitations of our work in this paper. First, we ran our evaluations only on smaller versions of the GPT-4, Gemini 1.5, and OLMO 2 models. It remains an open question whether our pattern of results would be similar for larger versions of these models and other model families. Second, for BLEnD and the human evaluation, we ran our evaluations using prompts that were relevant to ten national countries and cultures; a more comprehensive study would require the use of a wider range of national and regional cultures. With the momentum in the community to create more culturally salient resources, a more comprehensive study in the future will help identify gaps and interventions for the majority world. Finally, all of our evaluations concerned solely English-language prompts and outputs. While we take it to be an important goal for generative language models that they be able to produce culturally-informed outputs about any culture in the world’s most widely-spoken language, the landscape of cultural awareness becomes much more nuanced when one considers the rich variegation that exists in phrasing and dialect across a wide range of languages. We leave it to future work to examine whether the strategies used here can be adapted to a multi-lingual context.
Ethical Considerations
As generative large language models are developed and deployed rapidly across the globe, it is important to reflect on how we can improve user experience at a similar pace. The promise of model utility for a myriad of tasks such as that of a writing assistant, remains unfulfilled if the model is not beneficial or usable for a vast majority. With this work, we attempted to begin adaptation of techniques in NLP to further the cause of cultural awareness and relevance in models. As noted in our limitations, with more comprehensive work across a greater number of cultures and countries, we hope that development of more culturally aware models will be possible.
Appendix A Appendix
A.1 The Bespoke Knowledge Base

### v3

@@ Abstract
Generative large language models (LLMs) have demonstrated gaps in diverse cultural awareness across the globe. We investigate the effect of retrieval augmented generation and search-grounding techniques on LLMs’ ability to display familiarity with various national cultures. Specifically, we compare the performance of standard LLMs, LLMs augmented with retrievals from a bespoke knowledge base (i.e., KB grounding), and LLMs augmented with retrievals from a web search (i.e., search grounding) on multiple cultural awareness benchmarks. We find that search grounding significantly improves the LLM performance on multiple-choice benchmarks that test propositional knowledge (e.g., cultural norms, artifacts, and institutions), while KB grounding’s effectiveness is limited by inadequate knowledge base coverage and a suboptimal retriever. However, search grounding also increases the risk of stereotypical judgments by language models and fails to improve evaluators’ judgments of cultural familiarity in a human evaluation with adequate statistical power. These results highlight the distinction between propositional cultural knowledge and open-ended cultural fluency when it comes to evaluating LLMs’ cultural awareness.
@@ 1 Introduction
Contemporary large language models (LLMs) are pre-trained on huge corpora of natural language text Radford et al. (2019) and then fine-tuned using human feedback to improve their quality Bai et al. (2022). equal contribution During both processes, it is possible for text from a particular culture or cultures to be over-represented in the training data Dodge et al. (2021) and for the perspectives, norms, and mores of specific cultures to be over-represented in the feedback from human evaluators Prabhakaran et al. (2022); Atari et al. (2023). Consequently, there is growing recognition of generative LLMs’ shortcomings in representing and serving people from diverse geo-cultural backgrounds at the global scale Adilazuarda et al. (2024); Pawar et al. (2024); Agarwal et al. (2025). Models tend to stereotype different cultures Jha et al. (2023); Bhutani et al. (2024), erase and simplify their representation Qadri et al. (2025), and provide very limited knowledge and context about artifacts and norms that are salient to them Myung et al. (2024); Rao et al. (2024). Despite these gaps, strategies for eliciting culturally appropriate content from the models remain under-explored, with only some investigation of prompt engineering Rao et al. (2023); Wang et al. (2024) and model fine-tuning Chan et al. (2023); Li et al. (2024a, b) if not pre-training on more diverse non-English data.
@@
Here, we study two strategies to improve cultural awareness [^1] We use the term “cultural awareness” to refer to the general ability of LLMs to work well across cultures. It is conceptually similar to other terms in the literature, e.g., “cultural appropriateness” and “cultural informedness.” However, we are not attempting to define them precisely in this paper. of LLM generation using external knowledge. In the first strategy, we construct a bespoke cultural knowledge base (KB) and apply a retrieval augmented generation (RAG) technique Lewis et al. (2020); Gao et al. (2023) so that the input to the language model includes relevant cultural text from the knowledge base for better generation. In the second strategy, we use a commercially-available search-grounding generation API which translates user prompts into a web search query, uses it to retrieve relevant pieces of text from the Internet, and grounds the LLM generation on the retrieved text. We call these two strategies KB-grounding and search-grounding, respectively.
@@
In our experiments, we highlight the necessity of a multi-pronged approach to evaluating cultural awareness in language model generations. Specifically, we leveraged multiple benchmarks Myung et al. (2024); Rao et al. (2024); Bhutani et al. (2024) to evaluate cultural knowledge and the ability to avoid cultural stereotyping of the models equipped with the two strategies, compared with the vanilla generation baseline. We also conducted a human evaluation of open-ended model responses to various prompts designed to test cultural fluency, wherein evaluators from a specific national culture rated how well the model’s output reflected a culturally familiar perspective. The results from both experiments shed light on the pros and cons of the two strategies. Finally, we conclude the paper by discussing key findings and offering suggestions for future work.
2 Improving Cultural Awareness by Retrieving External Knowledge
@@
Retrieval augmented generation (RAG) is a technique for enhancing the quality of large language model generation. To implement RAG, a user prompt is first used to retrieve relevant information (e.g., text or documents) from a database, which is then added to the prompt before being passed to a generative LLM to produce a grounded response Lewis et al. (2020); Gao et al. (2023). RAG has shown to be effective in several applications, particularly those involving tasks that the base LLM was not well-trained for such as fact verification Asai et al. (2023); Singal et al. (2024); Khaliq et al. (2024) and domain-specific question answering (QA) Seo et al. (2024); Xiong et al. (2024); Kim and Min (2024). Some existing work also creates bespoke knowledge bases for use with RAG to tailor outputs to their specific applications Sun et al. (2025); Li et al. (2024c). Meanwhile, instead of querying a KB, other techniques retrieve external information by searching the internet, which allows access to up-to-date knowledge with high-quality ranking outputs Fan et al. (2024); Shuster et al. (2022); Lazaridou et al. (2022); Yao et al. (2022); Nakano et al. (2021); Komeili et al. (2022). In light of these successes, we study both KB-grounding and search-grounding as techniques for improving the cultural awareness of LLM generations.
@@
2.1 The Knowledge Base Grounding Strategy
We compiled culturally salient data from four large sources—CultureAtlas Fung et al. (2024), Cube Kannen et al. (2024), CultureBank Shi et al. (2024), and SeeGULL Jha et al. (2023)—to serve as our knowledge base as listed in Table 1. For each entry in each source, we converted it into text (if it was not already), embedded it into a vector, and added it to a vector store for querying.
Fig. 1 (top) shows how the KB-grounding strategy works. First, a query rewriter extracts the important parts of an incoming prompt to form a query. We can configure this step to test different KB queries (e.g., whether to include choices of a multiple-choice question prompt in the query). Next, we use the query to retrieve $n$ documents from the KB. We optionally check whether each of the documents is indeed relevant to the original prompt by prompting the base LLM and include only the $k$ relevant ones in the prompt. We call the process with the relevancy check step selective RAG, while non-selective RAG includes all $n$ documents in the prompt. Then, we feed the augmented prompt to the LLM to get a raw answer. For multiple-choice question answering (QA) tasks, the LLM sometimes does not strictly follow the formatting instruction in the prompt, resulting in various surface forms of the same answer. For example, the raw outputs for the answer “1) Yes” we found include, e.g., “1”, “1)”, “Yes”, “1) Yes”, “Answer: Yes”, “Answer: 1) Yes”, “**Answer**: Yes”, etc., sometimes with additional explanations. Therefore, we create a method, called a manual verbalizer, to normalize raw answers and map them to one of the choices so we can compute the model accuracy (similar to verbalizers in LLM-based text classification Schick and Schütze (2021); Thaminkaew et al. (2024)). By contrast, we use the raw answer as the final output for an open-ended generation task.
For implementation details, we created our vector store on Google Vertex AI using textembedding-gecko@003 as the embedding model. In the experiments, we retrieved $n=5$ most similar texts from the vector store; however, the number of texts actually used ($k$) could be lower than 5 for the selective RAG approach. In Section 3, we applied KB-grounding to three LLMs: gemini-flash-1.5 (Gemini) Team et al. (2024), gpt-4o-mini (GPT) OpenAI et al. (2024) and olmo2-7b (OLMo) OLMo et al. (2025), all with the temperature of 0.5. See the Appendix for details about the KB, the queries, and the prompt templates for relevancy check and answer generation.
2.2 The Search Grounding Strategy
Fig 1 (bottom) outlines how the search-grounding strategy works. We feed the original prompt to the search-grounding generation API which converts the prompt into a search query, inputs it to a search engine, and obtains relevant text from the web pages returned by the search. Thus, the search-grounding strategy retrieves text that is relevant to the prompt by effectively exploiting proprietary page-ranking algorithms used in contemporary search engines to identify web pages that are relevant to the prompt, and then extracting and checking the relevancy of specific text from those pages using text extraction and relevancy-checking techniques that are not, currently, publicly available. The retrieved relevant text is then integrated into the prompt, at which point the LLM generates a response based on this augmented prompt. Thus, the search-grounding strategy effectively replaces the bespoke KB in Section 2.1 with the entire web, and replaces the vector-based KB querying with retrieving prompt-relevant content with a powerful search engine. We implemented search grounding on Google Vertex AI. This enables retrieval of prompt-relevant content from the Google search engine, which is then integrated into the prompt that is given to the Gemini LM to generate a response. Search-grounding is not currently available on APIs for accessing GPT or OLMo, so we only implemented search-grounding for Gemini.
3 Cultural Competence Benchmarks
We evaluated the KB-grounding and search-grounding strategies on multiple-choice cultural QA benchmarks. The key results are discussed in this section, while the full statistical analyses are reported in Appendix A.5.
3.1 Cultural Knowledge
Setup.
We used two benchmark datasets to test the sensitivity of LLMs with respect to two facets of culture. The first, BLEnD, contains questions about everyday cultural knowledge (such as food, sports, family, and education) in different countries Myung et al. (2024). We used the ~24k English questions from BLEnD that are related to ten countries [^2] China, Ethiopia, Greece, Indonesia, Iran, Mexico, South Korea, Spain, the United Kingdom, and the United States. represented in our bespoke KB. The second, NormAd, focuses on cultural norms and values Rao et al. (2024). Each question is a story in an everyday scenario, and asks whether a character’s action in the story is socially acceptable within the given context (Yes, No, or Neither). We experimented with two types of contexts – Country and Country+Value. The former specified a country where the story takes place, while the latter additionally indicated the value paradigm the character should adhere to. NormAd has ~2.6k questions in total.
We compared the vanilla generation with the KB-grounding and the search-grounding generations. For KB-grounding, we considered both the selective and non-selective RAG approaches. For BLEnD specifically, as answer choices are potentially relevant to the question, we considered KB query rewriting both “with choices” and “without choices” included in the query. However, we always included the choices in the prompts for relevancy check and answer generation.
Results.
Fig. 2 (left) presents the results of the BLEnD benchmark. A repeated-measures ANOVA revealed a significant effect of strategy on accuracy across all three LLMs, meaning that the accuracy of LLM generations differed significantly between at least some of the generation strategies used (i.e., search-grounding, the varieties of KB-grounding, or vanilla). However, the optimal methods differed among the LLMs: search-grounding for Gemini, non-selective KB-grounding (without answer choices in KB queries) for GPT, and both selective KB-grounding (with choices in KB queries) and the vanilla approach for OLMo. Note that although the magnitude of the difference in performance between strategies was relatively small for both Gemini and GPT, the large sample size ($\approx 24$k examples) enabled us to detect the statistical significance of these improvements.
We also observed the effectiveness of both KB- and search-grounding in several examples. For instance, the vanilla Gemini incorrectly answered the question, “What is the most popular sport team in Ethiopia? (A) coffee (B) lg twins (C) persepolis (D) real madrid.” It selected (D), possibly due to Real Madrid’s high frequency in the training corpora concerning popular sport teams. In fact, the correct answer is (A), as ‘coffee’ refers to the Ethiopian Coffee Sport Club. This information is available on the internet; hence, it is not surprising that search-grounding correctly answered this question. Notably, search-grounded Gemini achieves 74.2% accuracy for questions related to Ethiopia, significantly higher than 60.3% of the vanilla baseline and 62.9% of the best KB grounding setting. Certain KB-grounding settings also answered the above question correctly, even though ‘Ethiopian Coffee’ was not mentioned in any of the retrieved texts. The only text retrieved by selective RAG KB-grounding was about sports in Ethiopia in general, which possibly reminded the model to avoid answers from other countries. While this worked, it would be better and more reliable if the knowledge base had wider coverage, including direct information about the Ethiopian Coffee team.
Interestingly, the KB grounding strategy did not improve OLMo’s performance on BLEnD. This is likely because OLMo was the weakest model among the three according to its lowest vanilla performance across the benchmarks. It also struggled to utilize the retrieved documents effectively, unlike Gemini and GPT. This issue was prominent when we used the non-selective KB grounding strategy, where the relevancy check step was not applied and the model was presented with all the five retrieved texts, some of which were irrelevant and lengthy. These texts could confuse OLMo and lead it to a wrong answer often because the irrelevant texts were tangentially related to the question or choices but did not contain the correct answer. Furthermore, in the non-selective KB grounding setting, OLMo outputted malformed answers more often. BLEnD requires the answer in the JSON format. Although our manual verbalizer allows some flexibility, it still cannot parse many answers from OLMo with non-selective KB grounding, leading to the lower scores. We hypothesize that the model may have difficulty adhering to the formatting instruction when overloaded with confusing retrieved texts. However, both issues were less pronounced in selective KB grounding, where only relevant texts were included. Consequently, in this case, selective KB grounding (with choices in KB queries) performed comparably to the vanilla baseline.
The middle two charts in Fig. 2 present the results for NormAd for both Country and County+Value context types. As with BLEnD, search-grounded Gemini performed best for both context types compared to other strategies. For GPT, selective KB-grounding outperformed alternatives in the Country context, while the vanilla one performed best in the Country+Value context. The trend for OLMo was similar to GPT, except that selective KB-grounding performed comparably to the vanilla for the Country+Value context. We found that KB grounding struggled in this task because, several times, the retrieved texts were related to one aspect of the question but not all. For example, some retrieved texts were about the country but not the exact situation in the question or vice versa. This could cause the LLMs to give wrong answers. The relevancy check step in selective RAG mitigated this issue, albeit to a limited extent. We posit that a better-tuned KB retriever would further address this problem (if the pertinent fact is present in the KB Joren et al. (2024)).
3.2 Stereotype Avoidance
Setup.
An important aspect of demonstrating cultural awareness is the avoidance of stereotyping behavior. Hence, we adopted the SeeGULL stereotype avoidance test Bhutani et al. (2024) to evaluate the studied strategies. Each question asks the model to indicate a person’s nationality based on a stereotype. For example,
In every question, only one country in the choices is stereotypically associated with the question. However, we considered only ‘(?)’ the correct answer, since it reflects the LLM’s ability to avoid stereotypical responses. The benchmark contains 4.6k questions. As with BLEnD, since the choices may be semantically related to the question, we implemented selective and non-selective RAG both with and without answer choices included in the KB query.
Note that our bespoke KB and the resource used to construct this stereotype avoidance benchmark overlap as both rely on SeeGULL Jha et al. (2023), and this setup is intended. Unlike cultural knowledge evaluation in Section 3.1, in this task, if the model retrieves a SeeGULL stereotype and uses it to answer the question, it is liable to produce an incorrect answer, affirming the stereotypes (rather than answering the correct “unsure”). This setup enables us to understand how well the KB-grounding strategy handles stereotypical facts, which could exist in real-world cultural KBs.
Results.
In contrast to BLEnD, the non-selective KB grounding strategy significantly improved OLMo’s performance in this task because, when presented with several texts that seem irrelevant to the question, OLMo could not ground any choice with the retrieved texts and therefore answered “unsure”, which was considered a correct answer for this task. By contrast, Gemini may be able to ignore irrelevant texts and exploit the relevant ones, resulting in stereotypical answers.
To understand how stereotypical facts in the KB affected KB grounding, we examined the sources of the retrieved texts and found that, out of 4,600 questions, in the KB query “without choices” setting, 1,156 questions retrieved at least one SeeGULL stereotype. However, in only 35 of these did the stereotype’s content exactly match the question. Similarly, in the KB query “with choices” setting, 1,266 questions retrieved at least one SeeGULL stereotype, but the stereotype’s content exactly matched the question in only 2 cases. In both settings, SeeGULL stereotypes whose content matched the question always passed the relevancy check for GPT and OLMo and usually passed for Gemini ($\sim$80of the time). The inclusion of these stereotypes in the prompt flipped the original “unsure” answer of the vanilla model to a stereotypical answer. This suggests that including stereotypes in a prompt can induce a model to affirm stereotypes. However, as the KB has significantly fewer and narrower stereotypical contents than the internet, this issue is not as severe as the search-grounding strategy.
4 Human Evaluation
Setup.
To evaluate the strategies on a more open-ended text generation task, we translated five questions from BLEnD and five questions from NormAd into open-ended prompts asking the model to tell a story set in a particular country. We adapted each of these ten questions for the ten national cultures used in the BLEnD benchmark evaluation, leaving us with 100 (country, prompt) pairs. Next, we generated responses from Gemini using the vanilla, the selective KB-grounding, and the search-grounding. For each strategy, we generated three unique responses with the temperature of 0.5. Then we recruited nine evaluators from each of the ten studied cultures to rate, on a scale from 0 to 4, how culturally familiar each response was and to provide a brief justification of their score (see Appendices A.6A.7 for details).
Results.
[…]
5 Discussion and Conclusion
KB Grounding vs Search Grounding.
The advantages of search-grounding on BLEnD and NormAd speak to the vast space of cultural facts on the internet. Even the large KB we compiled here still lacked many culturally relevant facts. We also observed some bias in each knowledge source used: approximately 19% of CultureAtlas entries and 25% of CultureBank entries concern the culture of the United States. While the web as a whole remains biased towards Western sources and values Johnson et al. (2022), it is more likely to contain necessary cultural information due to its sheer scale. However, the poor performance of search-grounding on the stereotype avoidance benchmark reminds us that the context retrieved via web search could reinforce the (typically false) notion that stereotypes are factual and encourage models to affirm those stereotypes. This suggests that search-grounding is not yet a panacea for improving the cultural sensitivity of LLMs.
Knowledge vs Fluency.
The results of the human evaluation do not show that search-grounding or KB-grounding improved the cultural familiarity of LLM outputs. The divergence in the performance of both strategies on the human evaluation and the multiple-choice QA benchmarks suggests that we ought to draw a distinction between two varieties of cultural awareness. On the one hand, there is a sense of cultural awareness that involves possessing propositional knowledge about a culture (i.e., knowing facts about that culture), as measured by the multiple-choice QA benchmarks, where the grounding strategies can improve the model performance. However, another sense of cultural awareness involves writing and speaking like someone with first-hand experience of and immersion in a culture, i.e., a sense of cultural fluency, as measured by our human evaluation, which revealed that the two grounding strategies were of limited value. We leave it to future work to develop strategies for improving the cultural awareness of generative language models along this second axis.
Limitations
We acknowledge the several limitations of our work in this paper. First, we ran our evaluations only on smaller versions of the GPT-4, Gemini 1.5, and OLMO 2 models. It remains an open question whether our pattern of results would be similar for larger versions of these models and other model families.
Second, for BLEnD and the human evaluation, we ran our evaluations using prompts that were relevant to ten national countries and cultures; a more comprehensive study would require the use of a wider range of national and regional cultures. With the momentum in the community to create more culturally salient resources, a more comprehensive study in the future will help identify gaps and interventions for the majority world.
Third, we only implement the search-grounding strategy using the Gemini model, specifically using the “Grounding with Google Search” feature of Vertex AI. According to the metadata returned, this end-to-end API has its own methods for query rewriting, incorporating retrieved texts, and providing citations, the details of which are not publicly available. Since these steps are tied to Gemini, it is not replicable for other LLMs while maintaining a fair comparison. As more LLM APIs introduce search-grounding capabilities, we hope that the lessons learned from search-grounding Gemini, as presented in this paper, establish the necessity of carefully auditing the characteristics of any future search-grounded LLMs before using them in culturally sensitive applications. Moreover, the paper retains its core message that RAG (including search-grounding) can improve performance with respect to cultural propositional knowledge, but not necessarily cultural fluency.
Finally, all of our evaluations concerned solely English-language prompts and outputs. While we take it to be an important goal for generative language models that they be able to produce culturally-aware outputs about any culture in the world’s most widely-spoken language, the landscape of cultural awareness becomes much more nuanced when one considers the rich variegation that exists in phrasing and dialect across a wide range of languages. We leave it to future work to examine whether the strategies used here can be adapted to a multi-lingual context.
Ethical Considerations
As generative large language models are developed and deployed rapidly across the globe, it is important to reflect on how we can improve user experience at a similar pace. The promise of model utility for a myriad of tasks such as that of a writing assistant, remains unfulfilled if the model is not beneficial or usable for a vast majority. With this work, we attempted to begin adaptation of techniques in NLP to further the cause of cultural awareness and relevance in models. As noted in our limitations, with more comprehensive work across a greater number of cultures and countries, we hope that development of more culturally-aware models will be possible.
Appendix A Appendix
A.1 The Bespoke Knowledge Base